# LUMA: A Benchmark Dataset for Learning from Uncertain and Multimodal Data

### v1

@@ 4 Baseline models with Uncertainty Quantification
For uncertainty quantification, we implemented 3 approaches: Monte Carlo Dropout (MCD) [7], Deep Ensemble (DE) [21], Reliable Conflictive Multi-View Learning (RCML) [37]. In Monte Carlo Dropout and Deep Ensembles, we use the aleatoric entropy and the epistemic entropy as as uncertainties measures $H_{\text{Ale }}(y\mid\mathbf{x})=\operatorname{entropy}\left(p_{\text{Ale }}(y\mid\mathbf{x})\right)$ and $H_{\mathrm{Epi}}(y\mid\mathbf{x})=\operatorname{entropy}\left(p_{Epi}(y\mid\mathbf{x})\right)$, where $p_{Epi}$ and $p_{Ale}$ are the probabilities obtained according to [34]. In RCML, we measure the aleatoric uncertainty with the expected entropy such as:
$\mathbb{E}_{p(\boldsymbol{\pi}\mid\mathbf{x},\hat{\boldsymbol{\theta}})}[H[P(y\mid\boldsymbol{\pi})]]=-\sum_{k=1}^{K}\frac{\alpha_{k}}{\alpha_{0}}\left(\psi\left(\alpha_{k}+1\right)-\psi\left(\alpha_{0}+1\right)\right),$ (2)
@@
We evaluate the measures of accuracy and uncertainty of the models on the clean dataset, the dataset with reduced diversity ($\searrow$ Diversity), the dataset with increased sample noise ($\searrow$ Sample Noise), and the dataset with switched label $\nearrow$ Label Noise). The results are summarized in Table 1. (For noise generation parameters, please refer to the supplementary material). Since in RCML the uncertainty is quantified differently compared with MCD and DE, we cannot directly compare their values, and relative changes to the two type of uncertainty compared to the clean dataset is reported in the table.
@@ 5 Conclusion
While LUMA multimodal dataset and tools offer resources for effectively studying and extensively benchmarking multimodal uncertainty quantification, it’s important to acknowledge some limitations. 1) Potential bias in the text modality: The text data was generated by a large language model, which can introduce biases and errors. Despite manual and automatic efforts to mitigate such bias, the texts may still contain undetected biases or errors. Therefore, we emphasize that this dataset is strictly for research on multimodal learning and uncertainty quantification. 2) Limited audio samples: While all 42 in-distribution classes have at least 600 audio samples, our diversity reduction algorithm might not perform optimally for classes with only slightly more than 600 samples. This limitation could hinder studies focusing on the audio modality.
The dataset can be easily extended with additional modalities and augmented with more data samples. The open-source nature of the data compilation pipeline and code for uncertainty and noise generation facilitates the integration of new contributions from the community to promote multimodal uncertainty studies and benchmarking initatives.
Acknowledgments and Disclosure of Funding
We thank our colleagues and friends for contributing to the audio validation process, and to the HumanSignal team for providing an academic license for Label Studio.
This work is funded by the Institut Archimedes doctoral program of Aix-Marseille University (AMU).
Supplementary Material
In this section we present the supplementary material for LUMA dataset.
The supplementary material contains:
  • A - Details about Data Collection
  • B - Details about the included baseline architectures and datasets
  • C - Datasheet for LUMA dataset
Appendix A Details about Data Collection
A.1 Image Modality
A.1.1 Licenses of the source datasets
As mentioned in Section 3.2 of the main paper, we use the CIFAR-100 and CIFAR-10 [19] datasets and EDM [14] generated images from the same distribution for the image modality.
@@ A.2 Audio Modality
A.2.1 Licenses of the source datasets
As mentioned in Section 3.2 of the main paper, we collected audio data from 3 sources: The Spoken Wikipedia [16], LibriSpeech [28], and Mozilla Common Voice [2] corpora.
The Spoken Wikipedia dataset is published under the CC BY-SA 4.0 license. Following the requirements of this license, we also distribute our dataset under the same license. The audio samples extracted from this dataset are located in the https://huggingface.co/datasets/bezirganyan/LUMA/tree/main/audio/sw_audio directory in LUMA dataset.
The LibriSpeech dataset is published under the CC BY 4.0 license. The audio samples extracted from this dataset are located in the https://huggingface.co/datasets/bezirganyan/LUMA/tree/main/audio/ls_audio directory in LUMA dataset.
The Mozilla Common Voice dataset is published under the CC0 (public domain) license. The audio samples extracted from this dataset are located in the https://huggingface.co/datasets/bezirganyan/LUMA/tree/main/audio/cv_audio directory in LUMA dataset.
A.2.2 Audio extraction and validation
@@
Then, we perform automatic validation with the OpenAI’s Whisper Large V3 model [31], and on the subset of the negative predictions of the model we perform a manual validation. For the manual validation, we have collected manual annotations thanks to the efforts of our volunteering colleagues and friends. To build the annotation interface, we opted for the Label Studio [33], an open-source data labeling platform. You can see the screenshot of the annotation interface in Figure 6. In total, we collected 16,744 manual annotations from 17 human annotators over a period of 2 months. Each sample was annotated by 2 annotators, and only annotations with 100% agreement were accepted.
@@
A.3 Text Modality
For the text modality, as described in the Section 3.3, we employed Google’s Gemma-7B Instruct model [26] to generate more than 1,200 text samples per class, using 13 different prompts. The 13 prompts are as follows:
@@ A.3 Text Modality
[…]
% [TODO] Add audio generation pipeline
Appendix B Details about Baseline architectures and Datasets
@@ B.2 Baseline models
For the image modality, we used a simple convolutional neural network, as depicted in Figure 10. For the audio modality, we extracted 128x128 mel-spectrograms from padded audio samples, and used a convolutional network for classification, as depicted in Figure 1. For the text modality, we extracted the BERT [5] embeddings for each token, and averaged them out, so that we have one embedding per text passage. Then, we passed the embedding through a simple feed-forward neural network (Figure 12) to get the predictions. As depicted in Figures 10, 11 and 12, each model includes two output heads: one for the prediction and the other for aleatoric uncertainty, following the methodology outlined by Valdenegro-Toro and Mori [34].
@@
Then, to combine the aforementioned unimodal networks into a multimodal architecture, we adopted the late fusion approach. In the Monte Carlo Dropout and Deep Ensemble methods, we obtained the multimodal prediction by averaging the logits from the final layers of the classifiers. For the RCML, we modified the output of the last layer in each network to produce evidence, as described in [37], and followed their methodology for combining the evidence.
In Figures 10, 11 and 12 the Dropout layer is kept turned on during the inference stage, for Monte-Carlo Dropout method, and turned off for Deep Ensemble and RCML methods. The dropout probability for all models is 0.3. The number of networks in deep ensembles is 10.
@@
Appendix C Datasheet
In this section, we provide the datasheet [8] for LUMA dataset.
  1. 1.
    Motivation for Dataset Creation
    • Why was the dataset created? There is a lack of multimodal datasets that allow for controlled injection of different types of uncertainty for various data modalities. This dataset was created to address these needs and to provide a benchmark for learning from uncertain and multimodal data, offering a valuable resource to the research and R&D communities.
    • Has the dataset been used already? If so, where are the results so others can compare. We provide baseline results in this article, but the dataset has not been used yet in other research articles.
    • What (other) tasks could the dataset be used for? The dataset can be used to study the behavior of uncertainty quantification algorithms under different types of uncertainty in the data. The dataset can also be used for benchmarking multimodal deep learning models, robustness analysis, and to study modality imbalance [29] in multimodal learning.
    • Who funded dataset creation? This work is funded by Institut Archimedes doctoral program of Aix-Marseille University (AMU) in France.
  2. 2.
    Dataset Composition
    • What are the instances? In LUMA dataset, there are three modalities: image, audio, and text, belonging to 50 classes. Image modality contains images of different objects from CIFAR 10/100 classes [19] (e.g., horse, truck, man, bird, etc.). The audio modality contains audio tracks of people pronouncing these class labels. The text modality contains texts generated by Large Language Models (LLM) about the class labels.
    • How many instances are there in total (of each type, if appropriate)? In total, LUMA dataset has 135,127 audio samples, 62,875 text samples, 62,400 images from CIFAR-10/100, and around half a million generated images similar to CIFAR-100 data (belonging to the 50 classes). Nevertheless, LUMA dataset is a superset, and the open-source noise generation tool we provide takes a sample of 42 classes from each modality, with 500 training and 100 testing samples in each class. It also provides 3,859 Out-of-Distribution (OOD) samples.
    • What data does each instance consist of? Each instance consists of semantically related raw image, text and audio data.
    • Is there a label or target associated with each instance? Yes, each instance belongs one of 50 classes, which are selected from the CIFAR-10 and CIFAR-100 classes. The classes are: [‘man’, ‘boy’, ‘house’, ‘woman’, ‘girl’, ‘table’, ‘road’, ‘horse’, ‘dog’, ‘ship’, ‘bird’, ‘mountain’, ‘bed’, ‘train’, ‘bridge’, ‘fish’, ‘cloud’, ‘chair’, ‘cat’, ‘baby’, ‘castle’, ‘forest’, ‘television’, ‘bear’, ‘camel’, ‘sea’, ‘fox’, ‘plain’, ‘bus’, ’snake’, ‘lamp’, ‘clock’, ‘lion’, ‘tank’, ‘palm’, ‘rabbit’, ‘pine’, ‘cattle’, ‘oak’, ‘mouse’, ‘frog’, ‘ray’, ‘bicycle’, ‘truck’, ‘elephant’, ‘roses’, ‘wolf’, ‘telephone’, ‘bee’, ‘whale’].
    • Is any information missing from individual instances? There is no information missing in the LUMA dataset, but it is possible to reduce the information by the provided open-source noise generator tool to inject various types of uncertainty in the data.
    • Are relationships between individual instances made explicit? This is not applicable as we do not study the relationship between different data instances. Instead, we focus exclusively on the relationship between different modalities of the same instance, which are semantically related. For example, for the class cat, each instance contains an image of a cat, an audio pronouncing the word "cat", and a text passage about a cat.
    • Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set? The dataset is a superset containing all possible instances. The provided open-source tool shall be used to extract samples from the dataset with different amounts of uncertainties.
    • Are there recommended data splits (e.g., training, development/validation, testing)? If so, please provide a description of these splits, explaining the rationale behind them. Yes, the provided open-source tool extracts a recommended train and test sets with 600 and 100 samples respectively per class. For image data, we use the samples of the original train/test splits. For the other modalities, we split the data at random, trying to have a maximally diverse test set (see the main paper for the definition of diversity).
    • Are there any errors, sources of noise, or redundancies in the dataset? The text modality is generated through Gemma-7B Instruct [26] Large Language Model and can contain factual inaccuracies and biases. For the audio modality, we used automatic validation for 130,069 samples, as well as the help of 17 human annotators for manual validation of 8,372 samples.
    • Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)? The audio and text modalities are self-contained. For the image modality, due to the absence of an explicit license in the CIFAR dataset, we decided to provide the URL to the dataset and a code to download and align with the other modalities.
  3. 3.
    Collection Process
    • What mechanisms or procedures were used to collect the data? For the audio modality, three audio datasets (LibriSpeech [28], The Spoken Wikipedia [16] and Mozilla Common Voice [2]) were downloaded from their respective websites for later processing. For the image modality, CIFAR-10 and CIFAR-100 [19] datasets were downloaded from their website, and EDM [14] generated images were downloaded from https://huggingface.co/datasets/P2333/DM-Improves-AT. The text modality was generated using Gemma-7B Instruct [26] Large Language Model (LLM).
    • How was the data associated with each instance acquired? For the audio modality, each instance was obtained by extracting the pronunciation of specific words from audio/text parallel datasets using Algorithm 1. The audio samples were initially validated automatically using the Whisper speech-to-text model [31], followed by manual validation for the remaining samples. For the image modality, each instance was selected from a subset of images from the CIFAR-10 and CIFAR-100 datasets [19], as well as from images generated using the EDM generative model [14]. For the text modality, instances were generated using the Gemma-7B Instruct Large Language Model (LLM) [26] and automatically validated using the same language model.
    • If the dataset is a sample from a larger set, what was the sampling strategy The samples were selected to ensure an equal number of instances for each class. Additionally, based on user-specified parameters, the sampling method adjusts to include either highly diverse instances (uniform sampling) or less diverse ones, as detailed in Section 3.4.1.
    • Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)? Data collection was primarily carried out by the authors of this paper. For manual audio validation, our colleagues, including PhD and Master’s students as well as professors, volunteered to assist with the annotations. In total we had 16,744 annotations from 17 annotators.
    • Over what timeframe was the data collected? Does this timeframe match the creation timeframe of the data associated with the instances (e.g., recent crawl of old news articles)? If not, please describe the timeframe in which the data associated with the instances was created. The data was generated, collected, and validated between February 3, 2024, and May 5, 2024. The audio and image datasets were created much earlier. The Spoken Wikipedia dataset was released in 2017, the version of the Mozilla Common Voice dataset used was released on January 5, 2024, and the LibriSpeech dataset was released in 2015. The CIFAR datasets were released in 2009, and the EDM-generated images were released in February 2023.
  4. 4.
    Data Preprocessing
    • Was any preprocessing/cleaning/labeling of the data done? Yes, data preprocessing, cleaning and validation was performed for audio and text modalities, as described in Section 3.
    • Was the "raw" data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)? If so, please provide a link or other access point to the "raw" data. The "raw" data is not saved, but audio and image modalities are available online at the links provided in the main paper.
    • Is the software used to preprocess/clean/label the instances available? If so, please provide a link or other access point. The source code for data acquisition, pre-processing, cleaning and validation is not available. We do provide the source code for compiling the final versions of the LUMA dataset.
    • Does this dataset collection/processing procedure achieve the motivation for creating the dataset stated in the first section of this datasheet? If not, what are the limitations? Yes, the dataset allows to insert different types of noises to each of the modalities in a controlled manner.
  5. 5.
    Dataset Distribution
    • How will the dataset be distributed? (e.g., tarball on website, API, GitHub; does the data have a DOI and is it archived redundantly? The dataset is released in HuggingFace platform with DOI: 10.57967/hf/2502. The source code for LUMA uncertainty generation tool is open-sourced and available at https://github.com/bezirganyan/LUMA.
    • When will the dataset be released/first distributed? What license (if any) is it distributed under? The dataset is already distributed under the CC BY-SA 4.0 license.
    • Are there any copyrights on the data? The source audio datasets have CC BY-SA 4.0, CC BY 4.0 and CC0 licenses, and our use of the datasets corresponds to the terms in the licenses. The EDM Generated images have Apache-2.0 license. We could not find liceneses for the CIFAR datasets, hence, we merely provide links and source code to download these datasets from their official website.
    • Are there any fees or access/export restrictions? No
  6. 6.
    Dataset Maintenance
    • Who is supporting/hosting/maintaining the dataset? The dataset is hosted in the HuggingFace platform and will be maintained by the authors of the paper.
    • Will the dataset be updated? If so, how often and by whom? The dataset may be updated with more modalities by the authors of the paper. We will also accept contributions from the community.
    • How will updates be communicated? (e.g., mailing list, GitHub) The updated will be made explicit in the Dataset Card in the HuggingFace platform.
    • If the dataset becomes obsolete how will this be communicated? Any changes to the dataset or its relevance will be communicated in its HuggingFace page.
    • Is there a repository to link to any/all papers/systems that use this dataset? No
    • If others want to extend/augment/build on this dataset, is there a mechanism for them to do so? If so, is there a process for tracking/assessing the quality of those contributions. What is the process for communicating/distributing these contributions to users? We welcome contributions from the community. The contributors are welcome to open issues or pull requests in our GitHub and HuggingFace pages. If the contributions align with the purpose of the dataset, the authors will merge them into the respective repositories.
  7. 7.
    Legal and Ethical Considerations
    • Were any ethical review processes conducted (e.g., by an institutional review board)? No
    • Does the dataset contain data that might be considered confidential? No
    • Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety? Yes. The text modality was generated using the Gemma-7B instruct large language model, and hence, inherited the biases and stereotypes form the model. While we performed an automatic bias detection procedure as described in Section 3.3, there may still be samples that were not detected by our bias detection algorithms.
    • Does the dataset relate to people? No

### v2

@@ 4 Baseline models with Uncertainty Quantification
For uncertainty quantification, we implemented 3 approaches: Monte Carlo Dropout (MCD) (Gal & Ghahramani, 2016), Deep Ensemble (DE) (Lakshminarayanan et al., 2017), Reliable Conflictive Multi-View Learning (RCML) (Xu et al., 2024).
In Monte Carlo Dropout and Deep Ensembles, we use the aleatoric entropy and the epistemic entropy as as uncertainties measures $H_{\text{Ale }}(y\mid\mathbf{x})=\operatorname{entropy}\left(p_{\text{Ale }}(y\mid\mathbf{x})\right)$ and $H_{\mathrm{Epi}}(y\mid\mathbf{x})=\operatorname{entropy}\left(p_{Epi}(y\mid\mathbf{x})\right)$, where $p_{Epi}$ and $p_{Ale}$ are the probabilities obtained according to (Valdenegro-Toro & Mori, 2022). In RCML, we measure the aleatoric uncertainty with the expected entropy such as:
$\mathbb{E}_{p(\bm{\pi}\mid\mathbf{x},\hat{\bm{\theta}})}[H[P(y\mid\bm{\pi})]]=-\sum_{k=1}^{K}\frac{\alpha_{k}}{\alpha_{0}}\left(\psi\left(\alpha_{k}+1\right)-\psi\left(\alpha_{0}+1\right)\right),$ (2)
@@
We evaluate the measures of accuracy and uncertainty of the models on the clean dataset, the dataset with reduced diversity ($\searrow$ Diversity), the dataset with increased sample noise ($\searrow$ Sample Noise), and the dataset with switched label $\nearrow$ Label Noise). The results are summarized in Tables 1. (For noise generation parameters, please refer to the supplementary material). Since in RCML the uncertainty is quantified differently compared with MCD and DE, we cannot directly compare their values, and relative changes to the two type of uncertainty compared to the clean dataset is reported in the table. We observe that accuracy always decreases with increasing the label noise, but reducing diversity and increasing sample noise may not always decrease accuracy in the image modality.
@@
Figure 6: Changes in uncertainty estimations under different proportions of label noise (in percentages). The RCML (Reliable Conflictive Multi-View Averaging) approach consistently increases both aleatoric and epistemic uncertainties with increased label noise. In contrast, the MCD (Monte Carlo Dropout) and DE (Deep Ensemble) models sometimes fail to increase the corresponding uncertainty estimations in this experiment.
@@
To further evaluate the qualities of the uncertainties of the different models, we estimate the epistemic and aleatoric uncertainties under different amounts of label noise. Ideally, we expect a good uncertainty quantification algorithm to provide higher uncertainty values for more noisy data. As we can see from Figure 6, only RCML consitently raises the uncertainty estimates under increased label noise, which again shows the higher quality of its uncertainty estimates over the other baselines.
@@ 5 Conclusion
The dataset can be easily extended with additional modalities and augmented with more data samples. The open-source nature of the data compilation pipeline and code for uncertainty and noise generation facilitates the integration of new contributions from the community to promote multimodal uncertainty studies and benchmarking initiatives.
Appendices
Appendix A Dataset links and Licenses
In this section we present the dataset links and licenses for the LUMA dataset.
Appendix B Details about Data Collection
B.1 Image Modality
B.1.1 Licenses of the source datasets
As mentioned in Section 3.2 of the main paper, we use the CIFAR-100 and CIFAR-10 (Krizhevsky, ) datasets and EDM (Karras et al., 2022) generated images from the same distribution for the image modality.
@@ B.2 Audio Modality
B.2.1 Licenses of the source datasets
As mentioned in Section 3.2 of the main paper, we collected audio data from 3 sources: The Spoken Wikipedia (Köhn et al., 2016), LibriSpeech (Panayotov et al., 2015), and Mozilla Common Voice (Ardila et al., 2020) corpora.
The Spoken Wikipedia dataset is published under the CC BY-SA 4.0 license. Following the requirements of this license, we also distribute our dataset under the same license. The audio samples extracted from this dataset are located in the sw_audio directory in LUMA dataset.
The LibriSpeech dataset is published under the CC BY 4.0 license. The audio samples extracted from this dataset are located in the ls_audio directory in LUMA dataset.
The Mozilla Common Voice dataset is published under the CC0 (public domain) license. The audio samples extracted from this dataset are located in the cv_audio directory in LUMA dataset.
B.2.2 Audio extraction and validation
@@
Then, we perform automatic validation with the OpenAI’s Whisper Large V3 model (Radford et al., 2023), and on the subset of the negative predictions of the model we perform a manual validation. For the manual validation, we have collected manual annotations thanks to the efforts of our volunteering colleagues and friends. To build the annotation interface, we opted for the Label Studio (Tkachenko et al., 2020-2022), an open-source data labeling platform. You can see the screenshot of the annotation interface in Figure 7. In total, we collected 16,744 manual annotations from 17 human annotators over a period of 2 months. Each sample was annotated by 2 annotators, and only annotations with 100% agreement were accepted.
@@
B.3 Text Modality
For the text modality, as described in the Section 3.3, we employed Google’s Gemma-7B Instruct model (Mesnard et al., 2024) to generate more than 1,200 text samples per class, using 13 different prompts. The 13 prompts are as follows:
@@ B.3 Text Modality
Appendix C Details about Baseline architectures and Datasets
@@ C.2 Baseline models
For the image modality, we used a simple convolutional neural network, as depicted in Figure 11. For the audio modality, we extracted 128x128 mel-spectrograms from padded audio samples, and used a convolutional network for classification, as depicted in Figure 1. For the text modality, we extracted the BERT (Devlin et al., 2019a) embeddings for each token, and averaged them out, so that we have one embedding per text passage. Then, we passed the embedding through a simple feed-forward neural network (Figure 13) to get the predictions. As depicted in Figures 11, 12 and 13, each model includes two output heads: one for the prediction and the other for aleatoric uncertainty, following the methodology outlined by Valdenegro-Toro & Mori (2022).
@@
Then, to combine the aforementioned unimodal networks into a multimodal architecture, we adopted the late fusion approach. In the Monte Carlo Dropout and Deep Ensemble methods, we obtained the multimodal prediction by averaging the logits from the final layers of the classifiers. For the RCML, we modified the output of the last layer in each network to produce evidence, as described in (Xu et al., 2024), and followed their methodology for combining the evidence.
In Figures 11, 12 and 13 the Dropout layer is kept turned on during the inference stage, for Monte-Carlo Dropout method, and turned off for Deep Ensemble and RCML methods. The dropout probability for all models is 0.3. The number of networks in deep ensembles is 10.
@@
C.3 Classification accuracies for baseline models
In Table 4 we present the classification accuracy measures for the clean dataset and variations in accuracy under different types of noise for the three baseline models (Monte Carlo Dropout, Deep Ensemble, and RCML) across each modality. Each version of the noisy dataset incorporates a single type of noise: Reduced Diversity (K=20), Label Noise (30%), and Sample Noise, with severity varying by noise type and modality (detailed in supplementary material). The RCML model, a multimodal approach, is evaluated exclusively in the multimodal setting. We observe that accuracy always decreases with increasing the label noise, but reducing diversity and increasing sample noise may not always decrease accuracy in the image modality.
Table 4: Classification accuracies for the clean dataset and variations in accuracy under different types of noise.
Model Clean Dataset Reduced Diversity Increased Label Noise Increased Sample Noise
Accuracy Difference in accuracy from the clean dataset
MCD Image 0.335 +0.058 -0.306 +0.019
MCD Audio 0.867 -0.025 -0.784 -0.155
MCD Text 0.965 -0.027 -0.864 -0.144
MCD Multi. 0.991 -0.010 -0.874 -0.063
DE Image 0.387 +0.066 -0.166 +0.019
DE Audio 0.912 -0.003 -0.809 -0.149
DE Text 0.973 -0.023 -0.864 -0.125
DE Multi. 0.996 -0.006 -0.849 -0.042
RCML Multi 0.973 -0.128 -0.833 -0.148